# Making school choice lotteries transparent

### v2

@@ 1 Introduction
Many cities worldwide use centralized school admission mechanisms and assign school priorities to students based on broad criteria, such as geographic proximity, demographic characteristics, and other factors. [^1] For instance, the Boston public school system (https://www.bostonpublicschools.org) classifies students into categories such as present school priority, sibling priority, EEC/ELC priority, and East Boston/Non-East Boston priority. Public schools in New York City (NYC; https://www.schools.nyc.gov) rank students based on zoned priority, sibling priority, and district priority. Priority ties frequently occur within these systems (Erdil and Ergin, 2008; Abdulkadiroğlu, Pathak and Roth, 2009). Popular mechanisms, which require strict priority orders for input, often resort to lotteries to resolve these ties to address fairness concerns. Despite the crucial role that lotteries play in determining student assignments, in current practices, students are typically unaware of their lottery outcomes when submitting preferences. To our knowledge, this issue has been overlooked in the school choice literature, presumably for two reasons. First, many studies assume strict priorities from the outset, thereby eliminating the need to consider lotteries. Second, most studies assume that students can submit complete preference lists, eliminating the need for lottery information under strategy-proof mechanisms (e.g., deferred acceptance). However, in practice, preference lists are often constrained in length. In such circumstances, no mechanisms can be truly strategy-proof, as students must strategically choose which schools to rank (rather than how to rank the complete list of schools), resulting in the thorny issue of going unmatched. In NYC, while students can rank up to twelve schools, 7% (around 5,000 students) remained unmatched after the main round of the match during the years 2021 and 2022. In certain districts, such as Manhattan district (D2), the unmatch rate can reach as high as 18%. Consequently, the unmatch rate emerges as a metric that policymakers aim to reduce.
@@
This paper addresses the aforementioned issue in practical school choice. We advocate for a “reveal” policy in which lotteries are drawn and revealed to students before their submission of preferences, contrasting it with the current “cover” policy, whereby students are uninformed about their lotteries before submitting preferences. To illustrate the different consequences of the two policies, consider two students $i$ and $j$ competing for two schools A and B, each with a single seat. Both students prefer A over B, and they are in priority ties for both schools. A single lottery is used to break the ties. Students are limited to reporting only one school. Under the cover policy, provided that both students value A sufficiently higher than B, both will report A. This is a rational choice, despite their full awareness of the risk of being unmatched. In contrast, under the reveal policy, students can effectively coordinate their strategies upon learning their lottery numbers: the student favored by the lottery outcome will report A and the other will report B. Consequently, both students will be matched, and ex ante each still has an equal chance of being admitted to A.
The intuition behind the benefits of the reveal policy is simple. We demonstrate this intuition both theoretically and experimentally. In two variants of the canonical model of Abdulkadiroğlu, Che and Yasuda (2011) (ACY for short), we show that the reveal policy can effectively resolve uncertainties for students and simplify their strategies by informing them of the “right” schools to rank. Especially, the reveal policy can achieve the same effect as allowing students to report complete and true preferences. We quantify the benefits of the reveal policy in laboratory experiments.
@@
In ACY’s model, students have common ordinal preferences over schools but differ in their privately known cardinal utilities. Schools do not assign priorities and instead use a single lottery to rank students. The deferred acceptance (DA) mechanism determines the matching. This model is particularly well-suited for our study, as its simple priority structure and the conflicting preferences setting together amplify the role of lotteries.
@@
Our first model deviates from ACY by restricting the number of schools that students can submit. We show that a simple reveal policy, termed Reveal policy, in which students are informed of their own lottery numbers but not the others’, [^2] We assume that the number of students and the capacities of schools in our model are common knowledge. is sufficient to resolve uncertainties for students: each student ranks and is assigned to their best achievable school, resulting in a matching equivalent to that produced by DA with complete preference lists. Consequently, every student is always matched. From both the interim (after utilities are realized but before the lottery is drawn) and the ex-ante (before utilities and the lottery are drawn) perspective, students have equal chances of attending each school. In contrast, under the Cover policy (in which students cannot see their lottery numbers), students’ strategies can be complicated. For certain realized utilities, students may concentrate their submissions on a subset of schools, resulting in wasted seats at other schools and leaving some students unmatched. From the interim perspective, every student faces a positive probability of being unmatched, and from the ex-ante perspective, students receive a random assignment that is first-order stochastically dominated by that obtained under the Reveal policy.
@@
Our second model brings the analysis closer to real-world settings, where schools often implement nontrivial priority policies. To illustrate how the reveal policy can accommodate such priorities, we extend the first model by incorporating neighborhood priority: each school prioritizes a subset of students living in its neighborhood, while still resolving ties within the same group through a single lottery. The presence of neighborhood priority complicates the reveal policy. Consider a student who, upon learning his lottery number, applies to a school where he lacks neighborhood priority. In the first model, his lottery number directly indicates his ranking relative to all other students across all schools. However, with neighborhood priorities, he faces uncertainty about how many students with worse lottery numbers nonetheless hold neighborhood priority at the school in question. In other words, his lottery number no longer provides a clear signal of his chances at each school. In practice, however, this issue may be solved in several ways. For instance, if the student population is large and their preference distribution is predictable, the uncertainty diminishes due to the law of large numbers. Moreover, if students’ preference distribution remains relatively stable over time, the cutoff lottery numbers for schools will tend to stabilize across years. [^3] While this argument appears to be new in the context of lottery-based school choice, it has been empirically verified in test-based admission systems in countries such as China and Turkey. The underlying theoretical insight is provided by Azevedo and Leshno (2016), who show that a continuum market generically has a unique stable matching, which is the limit of stable matching in large finite markets. In such settings, students can use past cutoff data along with their received lottery number to estimate their chances at different schools with relative precision.
@@
Another approach policymakers may adopt is to provide additional statistics about lottery outcomes, alongside each student’s individual lottery number. To illustrate this idea, our second model considers a more informative version of the reveal policy, termed RevealMore policy, in which policymakers publicly disclose, for each neighborhood, the number of students who receive lottery numbers weakly above each possible value. [^4] Similar policies are used in China’s college admissions system, where students are privately informed of their exam scores while the government publishes the number of students scoring above each possible value. These statistics allow students to infer the distribution of lottery numbers across neighborhoods. We show that the RevealMore policy can replicate the effect of the Reveal policy in the first model by effectively informing students of their achievable schools. In contrast, under the Cover policy, students with certain utilities choose to remain in their neighborhoods to minimize risk, which undermines the core objective of school choice. [^5] The experiment of Calsamiglia, Haeringer and Klijn (2010) provided evidence that under the Cover policy, compared to unconstrained preference lists, constrained preference lists cause students to exhibit a greater neighborhood school bias and increased segregation.
@@
In both models, student strategies under the reveal policies are characterized straightforwardly, whereas their strategies under the cover policy are far less transparent. This contrast is not unique to our setting. Prior literature has shown that student strategies under DA with constrained preference lists are generally difficult to characterize, even under complete information and strict priorities (Haeringer and Klijn, 2009; Decerf and Van der Linden, 2021). This might explain why much of the literature focuses on complete preference lists, despite the departure from reality, and underscores the practical challenge facing students. Only recently have Ali and Shorrer (2025) provided a framework for understanding the optimal portfolio of schools students should submit when school admissions are correlated. They show that the portfolios students choose depend on their subjective beliefs. In school choice, a single tie-breaking lottery creates correlated admissions. It is conceivable that more pessimistic students might apply less aggressively, leading to less favorable outcomes. In this sense, the reveal policy may help level the playing field by eliminating disparities arising from differences in subjective beliefs, even though this potential benefit falls outside the scope of our formal analysis.
@@
To test our models and quantify the benefits of the reveal policy, we conduct two laboratory experiments. The main experiment adopts a relatively simple environment in which students rank a single school in their preference lists. Across treatments, we vary the presence of neighborhood schools and the lottery revelation policy. The simple environment allows us to characterize students’ equilibrium strategies in all treatments contingent on their roles and utility types. We derive precise predictions regarding the expected match rates and expected student payoffs for each treatment. The experimental results indicate that the Reveal policy outperform the Cover policy by yielding significantly higher match rates and student payoffs, regardless of the presence of neighborhood schools. Further, the Reveal policy significantly increases the likelihood that students with various lottery numbers adhere to equilibrium strategies. Notably, when neighborhood schools are present, the RevealMore policy enhances market efficiency and reduces deviations from equilibrium strategies, as predicted, compared to the Reveal policy. The robustness experiment implements a more complex environment with a larger market, permitting students to rank up to two schools in their preference lists. The superiority of the reveal policy persists across different strategic environments.
While our paper operates under the assumption of a single tie-breaking lottery, the main insight carries over to multiple lotteries (i.e., different schools may use different lotteries). However, in the context of multiple lotteries, students must navigate more complex information under the reveal policy, which leads us to believe that a single lottery is superior in this regard. Fortunately, single lotteries are more frequently employed in practice (e.g., in NYC). [^6] There is a body of literature that compares single lotteries versus multiple lotteries for tie-breaking (e.g., Abdulkadiroğlu, Pathak and Roth, 2009; Pathak and Sethuraman, 2011; Ashlagi and Nikzad, 2020; Arnosti, 2023).
@@
Although our models build on ACY, our insights differ significantly. ACY compare the Boston mechanism (BM) and DA under the cover policy with unconstrained preference lists. Their main message is that students’ strategic behavior under BM, in response to the uncertainties associated with tie-breaking, can result in higher welfare than DA. [^7] However, the experiments of Featherstone and Niederle (2016) suggest that real-world students are unlikely to coordinate on the non-truth-telling equilibrium of BM suggested by ACY. In contrast, we compare two lottery revelation policies under DA with constrained preference lists. The benefit of the reveal policy comes from the simplified strategies of students after learning their lottery results.
The reveal policy examined in this study resembles the reform of the preference submission timing in China’s college admission system (Lien, Zheng and Zhong, 2016, 2017). This reform shifts the timing of preference submission from before students know their college entrance exam scores to after they receive them. This reform reduces uncertainty for students and improves match quality. In school choice, lottery numbers play a role similar to exam scores but are exogenous and independent of students’ efforts. In China, even when students could not observe their exact scores, they are often able to estimate them with considerable accuracy. In contrast, students have no basis for predicting their lottery numbers in school choice. Therefore, we believe that the reveal policy can have a greater impact on student strategies than the reform in China.
Finally, it is notable that during the preparation of this paper, NYC adopted the reveal policy for the 2022-2023 season. Initially, NYC refused to reveal lottery numbers. However, following a parent-led campaign under the New York State’s Freedom of Information Law, NYC first agreed to reveal lotteries upon request after admissions and finally decided to disclose lotteries before applications. The empirical implications of this policy shift represent an exciting area for future research. NYC continues to progress toward providing more lottery information. For instance, it still withholds historical cutoff data. Using a crowdsourced survey in NYC, Marian (2023) estimates school cutoffs and demonstrates their positive impact on reducing the unmatch rate in subsequent years among the survey participants. Her analysis aligns with our insight in this paper.

### v3

@@ 1 Introduction
People need information to make informed decisions. We apply this logic to guide policy reforms in school choice. Many cities worldwide use centralized school admission mechanisms and assign school priorities to students based on broad criteria (geographic proximity, demographic characteristics, and other factors). Priority ties frequently occur within these systems. Popular mechanisms, such as the Boston and the Deferred Acceptance (DA), however, require strict priorities for input. So, to guarantee fairness, those cities often resort to lotteries to resolve priority ties. Despite the crucial role that lotteries play in determining student assignments, in current practices, students are typically unaware of their lottery outcomes when submitting preferences. To our knowledge, this issue has been overlooked in the school choice literature, presumably for two reasons. First, many studies assume strict priorities from the outset, thereby eliminating the need to consider lotteries. Second, most studies assume complete preference lists and advocate for strategy-proof mechanisms, under which students have no need to acquire any information. However, in practice, preference lists are frequently constrained in length. [^1] Additionally, forming a complete ranking of schools is itself often challenging: acquiring meaningful information about each school can be costly, time-consuming, and unevenly accessible across students. As a result, preference lists are not only administratively capped but also de facto constrained by these information-acquisition frictions. We will return to this point in the later part of this section. According to neilson2024rise, who reviews primary, secondary, and higher education markets in 149 countries, Chile is the only country that employs a centralized strategy-proof mechanism and an unlimited preference list in primary education; only a few other countries do so for secondary and higher education. Under constrained preference lists, no mechanisms can be truly strategy-proof, as students must strategically choose which schools to rank, and the risk of going unmatched becomes a nonnegligible issue. In New York City (NYC), while students can rank up to 12 schools, 7% (around 5,000 students) remained unmatched after the main round of the match during the years 2021 and 2022. In certain districts, the unmatch rate can reach as high as 18%.
@@
This paper addresses the above issue in practical school choice. We advocate for a “reveal” policy in which lotteries are drawn and revealed to students before their submission of preferences, contrasting it with the current “cover” policy, whereby students are uninformed about their lotteries before submitting preferences. To illustrate the different consequences of the two policies, consider two students $i$ and $j$ competing for two schools A and B, each with a single seat. Both students prefer A over B, and they are in priority ties for both schools. A single lottery is used to break the ties. Students are limited to reporting only one school. Under the cover policy, provided that both students value A sufficiently higher than B, both will report A. This is a rational choice, despite their awareness of the risk of being unmatched. In contrast, under the reveal policy, students can effectively coordinate their strategies upon learning their lottery numbers: the student favored by the lottery outcome will report A and the other will report B. Consequently, both students will be matched, and ex ante, each still has an equal chance of being admitted to A.
The intuition for the reveal policy is simple: providing students with their lottery information enables more informed preference reporting. This paper employs stylized models and laboratory experiments to demonstrate this intuition. Notably, the reveal policy is not only conceptual; it has been implemented in practice. During the preparation of this paper, we observe that NYC has adopted this policy since the 2022-2023 season. Initially, NYC refused to reveal lottery numbers. However, following a parent-led campaign under the New York State’s Freedom of Information Law, NYC first agreed to reveal lotteries upon request after admissions and finally decided to disclose lotteries before preference submissions. Therefore, the results of this paper provide justification for NYC’s recent policy change.
@@
To demonstrate the benefits of the reveal policy transparently, we study two variants of the canonical model of abdulkadiroglu2011resolving (ACY hereafter) and show that the reveal policy can effectively resolve uncertainties for students and simplify their preference submission strategies by informing them of the “right” schools to rank. In ACY’s original model, students have common ordinal preferences over schools but differ in their privately known cardinal utilities. All students are in priority ties for all schools, and a single lottery determines the ranking of students for all schools. The DA mechanism determines the matching. This simple priority structure and the conflicting preferences setting together are particularly well-suited for demonstrating the role of lotteries.
@@
Our first model deviates from ACY’s by restricting the number of schools that students can rank. Thus, students need to decide which schools to rank. We show that a simple reveal policy, termed Reveal policy, in which students are informed of their own lottery numbers but not the others’, [^2] We assume that the number of students and the capacities of schools in our model are common knowledge. is sufficient to resolve uncertainties for students. Each student must rank and is assigned to his best achievable school, resulting in a matching equivalent to that produced by DA under complete preference lists. Consequently, every student is always matched. From both the interim (after utilities are realized but before the lottery is drawn) and the ex-ante (before utilities and the lottery are drawn) perspective, students have equal chances of attending each school. In contrast, under the Cover policy, students’ strategies can be complicated. For certain realized utilities, students may concentrate their submissions on a subset of schools, resulting in wasted seats at other schools and leaving some students unmatched. From the interim perspective, every student faces a positive probability of being unmatched, and from the ex-ante perspective, students receive a random assignment that is first-order stochastically dominated by that obtained under the Reveal policy.
@@
Our second model brings the analysis closer to real-world settings, where schools often implement nontrivial priority policies. A frequently observed priority policy in the real world is neighborhood. So, we extend the first model by incorporating neighborhoods: each school prioritizes a subset of students living in its neighborhood, while still resolving ties within the same student type by a single lottery. The presence of neighborhoods complicates the reveal policy. Consider a student who, upon learning his lottery number, applies to a school where he lacks neighborhood priority. In the first model, his lottery number directly indicates his ranking relative to all others across all schools. In the second model, however, he faces uncertainty about how many students with worse lottery numbers nonetheless hold neighborhood priority for the school. In other words, his lottery number no longer provides a clear signal of his chances at each school. In practice, this issue may be solved in several ways. For instance, if the student population is large and their preference distribution is predictable, the uncertainty diminishes due to the law of large numbers. Moreover, if students’ preference distribution remains relatively stable over time, the cutoff lottery numbers for schools in admission outcomes will tend to stabilize across years. [^3] While this argument appears to be new in the context of lottery-based school choice, it has been empirically verified in test-based admission systems (e.g., China and Turkey). The underlying theoretical insight has been provided by azevedo2016supply, who show that a continuum market generically has a unique stable matching, which is the limit of stable matchings in large finite markets. In such settings, students can use past cutoff data along with their lottery number to estimate their chances at different schools.
@@
Our second model considers another approach policymakers may adopt: providing students with additional statistics about lottery outcomes. To illustrate this idea, we propose a more informative version of the reveal policy, termed RevealMore policy, in which policymakers publicly disclose, for each neighborhood, the number of students who receive lottery numbers weakly above each possible value. [^4] Similar policies are used in China’s college admissions system, where students are privately informed of their exam scores while the government publishes the number of students scoring above each possible value. These statistics allow students to infer the distribution of lottery numbers across neighborhoods. We show that the RevealMore policy can replicate the effect of the Reveal policy in the first model by effectively informing students of their achievable schools. In contrast, under the Cover policy, students with certain utilities choose to remain in their neighborhoods to minimize risk, which undermines the objective of school choice. [^5] The experiment of calsamiglia2010constrained provided evidence that under the Cover policy, compared to unconstrained preference lists, constrained preference lists cause students to exhibit a greater neighborhood school bias and increased segregation. Interestingly, we observe that the RevealMore policy has been somehow adopted by NYC in the 2024-2025 season: in the application platform (MySchools), for each program, a student can see an icon next to the program indicating whether he has a low, medium, or high chance of getting accepted into the program, which is calculated based on last year’s admissions process, the student’s characteristics, and his assigned lottery number. [^6] Students will see one, two, or three orange bars next to each program. Three orange bars means that a student has a high chance of receiving an offer to this program. Two orange bars means a medium chance, and one orange bar means a low chance.
@@
In both models, students’ strategies under the reveal policies are straightforward, whereas their strategies under the cover policy are far less transparent. This is not surprising. Prior literature has shown that, even under strict priorities and complete information, students’ strategies under DA with constrained preference lists are difficult to characterize (haeringer2009constrained; decerf2021manipulability). Only recently have ali2025hedging provided a framework for understanding the portfolio of schools students should rank when school admissions are correlated. The portfolios students rank depend on their subjective beliefs. It is conceivable that more pessimistic students might apply less aggressively, leading to less favorable outcomes. Therefore, the reveal policy may help level the playing field by eliminating disparities arising from differences in subjective beliefs, though this potential benefit falls outside the scope of this paper.
@@
To test our models and quantify the differences between various policies, we conduct two laboratory experiments. The main experiment adopts a relatively simple environment in which we vary the presence of neighborhood schools and the lottery revelation policy, and students rank a single school in their preference lists. This one-school setting is extremely simple, but it is necessary to characterize students’ equilibrium strategies in all treatments. We derive precise predictions regarding the expected match rates and expected student payoffs. The experimental results indicate that the Reveal policy outperform the Cover policy by yielding significantly higher match rates and student payoffs, regardless of the presence of neighborhood schools. Further, the Reveal policy significantly increases the likelihood that students with various lottery numbers adhere to equilibrium strategies. Notably, when neighborhood schools are present, the RevealMore policy enhances market efficiency and reduces deviations from equilibrium strategies, as predicted, compared to the Reveal policy. The robustness experiment implements a more complex environment with a larger market and permits students to rank up to two schools. The superiority of the reveal policy persists across different strategic environments.
Overall, our results support the reveal policy and identify the key channel through which it improves the match process and outcome. To keep our analysis concise and focused, we make simplifying assumptions in models and experiments. In particular, the reveal policy in our models can eliminate all uncertainties for students. In practice, however, it is likely that the reveal policy can reduce but not entirely eliminate uncertainties. To what extent the policy can help students predict their achievable schools in practical markets is an empirical question. Since real-world markets are much larger and more complex than any laboratory experimental setting, this question has to be answered using field data. This represents an exciting direction for future research. Using a crowdsourced survey, marian2023algorithmic demonstrates the positive impact of NYC’s reveal policy on improving the match outcomes for survey participants. Her finding provides an empirical support for the insight in our paper.
Our analysis assumes that students face constrained preference lists, which holds widely in the real world. But since list length is itself a policy choice, this raises an immediate question: if policymakers seek to assist students, why not simply extend list lengths? More generally, how should we compare the effect of revealing lottery information with the effect of lengthening lists? The second question is again empirical, while the first connects to a deeper issue in school choice. Models since the seminal work of abdulkadiroglu2003school, including ours, typically assume that students have complete preferences before the match, but in reality students need information to form preferences and, under list constraints, must decide which schools to rank. NYC’s recent reform sheds light on this point. While NYC imposes the 12-school cap on preference list, many students already struggle to find 12 choices to list. In recent years, about 38% of high school applicants submitted full 12-school lists. For the remaining 62%, submitting shorter lists does not indicate that the length constraint was not binding for them; it reflects the difficulty in finding suitable schools to rank, in the absence of information. As of the 2024–2025 cycle, NYC has removed the 12-school cap on high school applications, allowing students to rank an unlimited number of schools. This change does not replace the reveal policy adopted in 2022–2023. Rather, NYC has increased the information provided to students (the icon described above) to help them evaluate admission chances at each school. This practice suggests that list length is not the only issue in practice. Providing information, which includes lottery results in the NYC case, remains essential for students to make informed decisions, even when lists are unconstrained.
This paper contributes to the growing literature on information provision in school choice. We discuss this literature below. Here, we highlight recent findings from empirical studies. Although Chile allows students to rank an unlimited number of schools, arteaga2022smart find that students on average rank only three schools, and this is because students lack information about which schools are in their choice set. Students typically overestimate their admission chances, and a significant portion of students face an over 50% probability of non-placement. In intervention experiments, students who initially submit short lists add one or more schools to their lists after being warned of the risk. Similarly, ainsworth2023households and ajayi2025information, using data from Romanian and Ghana, respectively, find that students often make school choice decisions with inadequate information, and, if appropriate information is provided, they respond sensitively and make better decisions. These findings suggest that researchers should not only work on the design of algorithms that produce matchings but also explore the design of information provision mechanisms. In our setting, since lottery results are free information on the platform, it could be provided at no cost to help students better estimate their choice set and secure better placements.
Related literature
To our knowledge, we are the first to formalize the reveal policy suggestion for lottery-based school choice. Importantly, it is not merely an idea but has already become practice in NYC, one of the world’s largest school choice systems. Previous studies that consider lotteries fall mainly into three categories: (1) examining how tie-breaking affects the efficiency of stable matching mechanisms; (2) reevaluating manipulable mechanisms by incorporating the uncertainty created by tie-breaking; (3) comparing different tie-breaking rules. All these studies take the prevailing practice of withholding lottery information from students as given.
Specifically, erdil2008s and abdulkadirouglu2009strategy show that the stability constraints created by exogenous tie-breaking can result in efficiency loss in the outcome of DA, and they propose algorithms to eliminate the efficiency loss. However, it is impossible to improve on the DA outcome without losing its strategy-proofness property.
@@
abdulkadiroglu2011resolving compare the Boston mechanism (BM) and DA under the cover policy with unconstrained preference lists. Their main message is that students’ strategic behavior under BM, in response to the uncertainties associated with tie-breaking, can result in higher welfare than DA. Our models build on ACY, but our insights differ significantly. We compare two revelation policies under DA with constrained preference lists. The benefit of the reveal policy comes from the simplified strategies of students after learning their lottery results, whereas the welfare superiority of BM in ACY’s result relies on non-truth-telling equilibrium, which, as suggested by the experiments of featherstone2016boston, might not be achievable by real-world students.
There is a body of literature that compares single lotteries versus multiple lotteries (i.e., different schools may use different lotteries) for tie-breaking (e.g., abdulkadirouglu2009strategy; pathak2011lotteries; ashlagi2020matters; arnosti2023lottery). While our paper operates under the assumption of a single lottery, the main insight carries over to multiple lotteries. However, under multiple lotteries, students must navigate more complex information, which leads us to believe that a single lottery is superior in this regard. Fortunately, single lotteries are more frequently employed in practice.
The reveal policy we study parallels the reform of preference-submission timing in China’s college admission system. The reform moves preference submission from before students learn their college-entrance exam scores to after scores are released. It is accompanied by a shift from sequential mechanisms (variants of BM) to parallel mechanisms (variants of DA), as studied by chen2017chinese. These changes are widely viewed as successful in reducing uncertainty for students and improve match quality. lien2016preference; lien2017ex, however, provide a different perspective on this reform. Assuming that the true abilities of students are not perfectly signaled by their exam scores, the authors highlight the superiority of BM with pre-exam preference submission. They argue that, when students submit their preferences before the exam, those students with higher abilities are more willing to apply for better schools since they have higher expected scores, while those with lower expected scores are less willing to do so. This produces a more ex-ante fair matching. Moreover, students having higher preference intensities for good schools are more willing to apply for them than those with lower preference intensities, which leads to a more ex-ante efficient matching. This latter point resembles ACY’s insight.
In school choice, lotteries play a role similar to exam scores, except that they are randomly generated and independent of students’ effort. The reform in China gives us confidence that a reveal policy in school choice can similarly help students. NYC’s adoption of this policy further suggests that policymakers have recognized its benefits. While lien2016preference; lien2017ex provide reflections on the reform in China, their arguments do not apply to school choice, since there is no concern about students’ abilities here, and since lottery numbers are random, the sorting channel central to their arguments is absent. Beyond conveying different messages, the models and experiments in lien2016preference; lien2017ex also differ from ours. They compare BM and Serial Dictatorship in one-to-one school matching with complete preferences lists. The difficulties associated with constrained preference lists are absent in their studies.
A growing literature studies how explicitly disclosing agents’ attainable choice sets affects their behavior and welfare in matching environments. A main insight is that sequentially disclosing information or presenting mechanisms in a dynamic form can improve information efficiency and simplify agents’ strategies. For example, li2017obvious demonstrates that dynamic mechanisms presenting agents with simple, staged choice sets can mitigate strategic mistakes made by boundedly rational agents. bo2020iterative; bo2024pick show that iterative versions of DA, which require students to apply to one school at a time, outperform the static DA in truth‐telling rates and in the frequency of stable outcomes. hakimov2023costly find that a sequential serial dictatorship mechanism induces less wasteful information acquisition and yields higher student welfare than its direct counterpart. [^7] Other contributions in this literature include bade2015serial, chen2021information; chen2022information, noda2022strategic, mackenzie2022menu, pycia2023theory, among others. In our setting, revealing the realization of the single tie-breaking lottery turns a complicated distribution over admissions into a simple, individualized choice set and thereby makes constrained rank-order list decisions more transparent and less risky.
Finally, our paper is related to the emerging literature on transparency in school choice and matching mechanisms. hakimov2025improving theoretically and experimentally study policies that allow students to verify their school assignments and the allocation rules under popular mechanisms. moller2025transparent formalizes a concept of transparency and examines conditions under which mechanisms promised by policymakers with limited commitment can be trusted or deviations can be detected. [^8] Similar concepts include credibility by akbarpour2020credible and auditability by grigoryan2024theory and pycia2024ordinal. While we do not study policymakers’ commitment problems per se, the reveal policy effectively works as a transparency intervention in our setting. Beyond improving efficiency and fairness, this intervention may bring an additional behavioral benefit not studied in our paper: the reduction in student anxiety. When students know that lotteries play a decisive role in their assignments yet remain unobservable, they may experience uncertainty and discomfort. Making lottery results visible can mitigate this psychological burden. Indeed, this concern appears salient in practice: in New York City, parents have organized campaigns calling for public disclosure of lottery results.